# Random Sperner lemma and random Brouwer fixed point theorem

### v3

@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
Theorem 3.5 below establishes a key connection between the set of proper $L^{0}$-labeling functions of $\mathscr{S}$ and the set of proper labeling functions of $\mathscr{S^{\prime}}$. In particular, (3) of Theorem 3.5 below gives the representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions, which considerably simplifies the proof of Theorem 2.9 below.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an $(n-1)$-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$ and $\mathscr{S}$ be an $L^{0}$-simplicial subdivision of $S$. Then we have the following statements:
@@
    For any proper $L^{0}$-labeling function $\phi$ of $\mathscr{S}$ and any $y\in ext(\mathscr{S}^{\prime})$, there exists a family $\{\psi_{i}:i\in\mathcal{N}_{\phi(y)}\}$ of proper labeling functions of $\mathscr{S^{\prime}}$ with $\psi_{i}(y)=i$ for each $i\in\mathcal{N}_{\phi(y)}$, namely
@@
(1). For any given $y\in ext(\mathscr{S}^{\prime})\subseteq S^{\prime}$, there exists $\{\alpha_{i},i=1\sim n\}\subseteq[0,1]$ with $\sum_{i=1}^{n}\alpha_{i}=1$ such that
@@
(2). For each $i\in\mathcal{N}_{\phi(y)}$, since $\mathcal{N}_{\phi(y)}\subseteq\chi(y)$, it is clear that there exists a proper labeling function $\psi_{i}$ of $\mathscr{S^{\prime}}$ such that $\psi_{i}(y)=i$, namely
@@
(3). For each $k\in\{1,2,\cdots,K\}$, let $y_{k}=\sum_{i=1}^{n}\lambda_{i}^{k}x_{i}$, where $\{\lambda_{i}^{k},i=1\sim n\}\subseteq[0,1]$ are such that $\sum_{i=1}^{n}\lambda_{i}^{k}=1$. Since $\phi$ maps $ext_{L^{0}}(\mathscr{S})$ into $L^{0}(\mathcal{F},\{1,2,\cdots,n\})$, for each $k\in\{1,2,\cdots,K\}$, there exists $\{a_{i}^{k},i=1\sim n\}\in p(1)$ such that
@@
of $\{a_{i}^{1},i=1\sim n\},\cdots,\{a_{i}^{K-1},i=1\sim n\}$ and $\{a_{i}^{K},i=1\sim n\}$, where $\{1,2,\cdots,n\}^{K}$ is the $K$-th Cartesian power of $\{1,2,\cdots,n\}$. Further, let
$I=\{(i_{1},i_{2},\cdots,i_{K})\in\{1,2,\cdots,n\}^{K}:a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}>0\},$
@@
$\mathcal{N}_{\phi(y_{k})}(\text{namely}~{}P_{k}(I))=\{p_{k,l}:1\leq l\leq M_{k}\},$
@@
that is to say, $I_{k,l}=P_{k}^{-1}(p_{k,l})$, then it is clear that $I=\bigcup_{l=1}^{M_{k}}I_{k,l}$ and $\{I_{k,l}:1\leq l\leq M_{k}\}$ is a pairwise disjoint family. Now, by (2) as has been proved, there exists a family $\{\varphi_{k,l}:1\leq l\leq M_{k}\}$ of proper labeling functions of $\mathscr{S}^{\prime}$ such that $\varphi_{k,l}(y_{k})=p_{k,l}$ for each $l\in\{1,2,\cdots M_{k}\}$, namely
@@
For each $k\in\{1,2,\cdots,K\}$ and each $l\in\{1,2,\cdots,M_{k}\}$, we define a proper labeling function $\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}$ of $\mathscr{S}^{\prime}$ for each $(i_{1},i_{2},\cdots,i_{K})\in I_{k,l}$ by $\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}=\varphi_{k,l}$, then one can easily see that
@@
Lemma 3.3, Lemma 3.4 and Theorem 3.5 show that $L^{0}$-simplexes, $L^{0}$-simplicial subdivisions and proper $L^{0}$-labeling functions are closely connected to their classical counterparts. These results motivate us to make full use of Lemma 2.8 to complete the proof of Theorem 2.9.
We can now prove Theorem 2.9.
@@
Further let $S_{m}=Conv_{L^{0}}(\{y_{1}^{m},\cdots,y_{n}^{m}\})$ for each $m\in\{1,2,\cdots,M\}$, then by (1) of Lemma 3.3,
@@
is an $(n-1)$-dimensional $L^{0}$-simplex contained in $S$. Furthermore, by (3.1), we have
@@
4 Proof of Theorem 2.1
@@
Definition 4.1 here is adopted from [19, 20] by following the traditional nomenclature of random metric spaces and random normed spaces (see [47, Chapters 9 and 15]).
@@
[…]
@@
Clearly, when $(\Omega,\mathcal{F},P)$ is trivial, namely =$\mathcal{F}=\{\Omega,\emptyset\}$, then an $RN$ module with base $(\Omega,\mathcal{F},P)$ just reduces to an ordinary normed space.
Let $(E,\|\cdot\|)$ be an $RN$ module $(E,\|\cdot\|)$ over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$. For any real numbers $\varepsilon>0$ and $10<\lambda<$ 1, let
@@
[…]
$B(x,r)=\{y\in E:\|y\|<r~{}\text{on}~{}\Omega\},$
[…]
@@
Define the $L^{0}$-inner product $\langle\cdot,\cdot\rangle:L^{0}(\mathcal{F},\mathbb{K}^{d})\times L^{0}(\mathcal{F},\mathbb{K}^{d})$ by
@@
where $\bar{y}_{i}$ denotes the complex conjugate of $y_{i}$ for each $i\in\{1,2,\cdots,n\}$. Then $(L^{0}(\mathcal{F},\mathbb{K}^{d}),\langle\cdot,\cdot\rangle)$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete random inner product module (briefly, a $\mathcal{T}_{\varepsilon,\lambda}$-complete $RIP$ module, see [22] for details). let $\|x\|=\sqrt{\langle x,x\rangle}$ for any $x\in L^{0}(\mathcal{F},\mathbb{K}^{d})$, then $(L^{0}(\mathcal{F},\mathbb{K}^{d}),\|\cdot\|)$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete $RN$ module.
@@
The randomized version of the Bolzano-Weierstrass theorem [31] states that for any a.s. bounded sequence $\{\xi^{0}_{n},n\in\mathbb{N}\}$ of random vectors from $(\Omega,\mathcal{F},P)$ to $\mathbb{R}^{d}$ (namely $\sup\{|\xi^{0}_{n}(\omega)|:n\in\mathbb{N}\}<+\infty$ $a.s.$) there exists a sequence $\{n_{k}^{0},k\in\mathbb{N}\}$ of positive integer-valued random variables on $(\Omega,\mathcal{F},P)$ such that $\{n_{k}^{0}(\omega),k\in\mathbb{N}\}$ is strictly increasing for each $\omega\in\Omega$ and $\{\xi^{0}_{n_{k}^{0}},k\in\mathbb{N}\}$ converges a.s., where $\xi^{0}_{n_{k}^{0}}(\omega)=\xi^{0}_{n_{k}^{0}(\omega)}(\omega)=\sum^{\infty}_{l=1}I_{(n_{k}^{0}=l)}(\omega)\xi^{0}_{l}(\omega)$ for any $k\in\mathbb{N}$ and any $\omega\in\Omega$ (as usual, $(n_{k}^{0}=l)$ denotes the set $\{\omega\in\Omega:n_{k}^{0}(\omega)=l\}$ for any $k$ and $l$ in $\mathbb{N}$). Motivated by this fundamental result, Guo et al. introduced the following notion of random sequential compactness in $RN$ modules.
@@
(1) In Definition 4.3, we require $n_{k}$ to be an element in $L^{0}(\mathcal{F},\mathbb{N})$, rather than a positive integer-valued random variable as in [26, Definition 2.10]. This choice is made for simplifying written expressions, and it is easy to check that the two formulations are essentially equivalent. (2) In the sense of Definition 4.3, the randomized Bolzano-Weierstrass theorem can be restated as: any a.s. bounded nonempty subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is random relatively sequentially compact. Moreover, as pointed out in [26], if $G$ is $\sigma$-stable, then $G$ is random relatively sequentially compact iff $\bar{G}_{\varepsilon,\lambda}$ (the closure of $G$ under $\mathcal{T}_{\varepsilon,\lambda}$) is random sequentially compact. Consequently, any a.s. bounded $\sigma$-stable nonempty $\mathcal{T}_{\varepsilon,\lambda}$-closed subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (for example, an $L^{0}$-simplex) must be random sequentially compact. (3) For any $(n-1)$-dimensional $L^{0}$-simplex $S$ of an $RN$ module, as pointed out by [26, Lemma 5.5], there exists an $L^{0}$-affine $\mathcal{T}_{\varepsilon,\lambda}$-isomorphism $T$ from $\Delta_{n-1}$ to $S$, where $\Delta_{n-1}:=Conv_{L^{0}}(\{e_{1},e_{2},\cdots,e_{n}\})$ is the standard $(n-1)$-dimensional $L^{0}$-simplex of $L^{0}(\mathcal{F},\mathbb{R}^{n})$. Since $\Delta_{n-1}$ is random sequentially compact as discussed above, and $T$ is $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous, then $S$ is random sequentially compact.
@@
    $a.s.$ sequentially continuous at $x_{0}\in G_{1}$ if $\{\|T(x_{n})-T(x_{0})\|_{2},n\in\mathbb{N}\}$ converges $a.s.$ to 00 for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G_{1}$ such that $\{\|x_{n}-x_{0}\|,n\in\mathbb{N}\}$ converges $a.s.$ to 00. Further, $T$ is said to be $a.s.$ sequentially continuous if $T$ is $a.s.$ sequentially continuous at any point in $G_{1}$.
@@
It is clear that a.s. sequential continuity implies $\mathcal{T}_{\varepsilon,\lambda}$-continuity. Moreover, when $G_{1}$ is $\sigma$-stable, the $\mathcal{T}_{\varepsilon,\lambda}$-continuity of a mapping $T$ from $G_{1}$ to $G_{2}$ necessarily implies random sequential continuity of $T$. Furthermore, [26, Lemma 4.3] shows that a $\sigma$-stable mapping $T$ of $G_{1}$ to $G_{2}$ is random sequentially continuous iff it is $\mathcal{T}_{c}$-continuous.
For any $(n-1)$-dimensional $L^{0}$-simplex $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ in an $RN$ module $E$ and any $L^{0}$-simplicial subdivision $\mathscr{S}=\{S_{i}:i\in I\}$ of $S$, the random diameter of $\mathscr{S}$ is defined by
@@
Let $(E,\|\cdot\|)$ be an $RN$ module over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$, $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ an $(n-1)$-dimensional $L^{0}$-simplex in $E$ and $\mathscr{S}$ an $L^{0}$-barycentric subdivision of $S$. Then
@@
First, we claim that $diam(G)=\bigvee_{k,l=1}^{m}\|y_{i}-y_{j}\|$ for any $(n-1)$-dimensional $L^{0}$-simplex $G=Conv_{L^{0}}(\{y_{1},y_{2},\cdots,y_{n}\})$ in $E$. In fact, it is clear that $\bigvee_{k,l=1}^{m}\|y_{k}-y_{l}\|\leq diam(G)$. On the other hand, for any $x=\sum_{i=1}^{m}\lambda_{i}y_{i}$ and $y=\sum_{i=1}^{m}\mu_{i}y_{i}$ in $G$, where $\lambda_{i},\mu_{i}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,m\}$ and $\sum_{i=1}^{m}\lambda_{i}=\sum_{i=1}^{m}\mu_{i}=1$, we have
@@
Second, for any $\pi\in G(n)$ and any $k,l\in\{1,2,\cdots,n\}$, without loss of generality, we can assume that $k>l$, then
@@
The key idea of the proof of classical Brouwer fixed point theorem for a continuous self-mapping on a simplex $S^{\prime}$ lies in constructing a proper labeling function $\psi$ on a given simplicial subdivision $\mathscr{S}^{\prime}$ of $S^{\prime}$ that is compatible with the continuous self-mapping on $S^{\prime}$, see Theorem 6.1 of [1] for details. Lemma 4.8 below provides a method for constructing such an $L^{0}$-labeling function in connection with an $(n-1)$-dimensional $L^{0}$-simplex $S$ in an $RN$ module, which improves the formulation of Lemma 2.2 of [11] in that the single proper $L^{0}$-labeling function provided here can be applied to the $L^{0}$-simplicial subdivision of every $(n-1)$-dimensional $L^{0}$-simplex contained in $S$.
@@
Let $(E,\|\cdot\|)$ be an $RN$ module with base $(\Omega,\mathcal{F},P)$, $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ an $(n-1)$-dimensional $L^{0}$-simplex in $E$ and $f:S\rightarrow S$ be a $\sigma$-stable mapping. Then there exists a mapping $\phi:S\rightarrow L^{0}(\mathcal{F},\{1,2,\cdots,n\})$ such that the following conditions are satisfied:

### v5

@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
Theorem 3.5 below establishes a key connection between the set of proper $L^{0}$-labeling functions of $\mathscr{S}$ and the set of proper labeling functions of $\mathscr{S^{\prime}}$. In particular, part (3) of Theorem 3.5 below gives the representation theorem of a proper $L^{0}$-labeling function by usual proper labeling functions, which considerably simplifies the proof of Theorem 2.9.
@@
Let $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ be an (n1)(n-1)-dimensional $L^{0}$-simplex in an $L^{0}$-module $E$ and $\mathscr{S}$ be an $L^{0}$-simplicial subdivision of $S$. Then we have the following statements:
@@
    For any proper $L^{0}$-labeling function $\phi$ of $\mathscr{S}$ and any $y\in ext(\mathscr{S}^{\prime})$, there exists a family $\{\psi_{i}:i\in\mathcal{N}_{\phi(y)}\}$ of proper labeling functions of $\mathscr{S^{\prime}}$ with $\psi_{i}(y)=i$ for each $i\in\mathcal{N}_{\phi(y)}$, namely,
@@
(1) For any given $y\in ext(\mathscr{S}^{\prime})\subseteq S^{\prime}$, there exists $\{\alpha_{i},i=1\sim n\}\subseteq[0,1]$ with $\sum_{i=1}^{n}\alpha_{i}=1$ such that
@@
(2) For each $i\in\mathcal{N}_{\phi(y)}$, since $\mathcal{N}_{\phi(y)}\subseteq\chi(y)$, it is clear that there exists a proper labeling function $\psi_{i}$ of $\mathscr{S^{\prime}}$ such that $\psi_{i}(y)=i$, namely,
@@
(3) For each $k\in\{1,2,\cdots,K\}$, let $y_{k}=\sum_{i=1}^{n}\lambda_{i}^{k}x_{i}$, where $\{\lambda_{i}^{k},i=1\sim n\}\subseteq[0,1]$ are such that $\sum_{i=1}^{n}\lambda_{i}^{k}=1$. Since $\phi$ maps $ext_{L^{0}}(\mathscr{S})$ into $L^{0}(\mathcal{F},\{1,2,\cdots,n\})$, for each $k\in\{1,2,\cdots,K\}$, there exists $\{a_{i}^{k},i=1\sim n\}\in p(1)$ such that
@@
of $\{a_{i}^{1},i=1\sim n\},\cdots,\{a_{i}^{K-1},i=1\sim n\}$ and $\{a_{i}^{K},i=1\sim n\}$, where $\{1,2,\cdots,n\}^{K}$ is the $K$-th Cartesian power set of $\{1,2,\cdots,n\}$. Further, let
$I=\{(i_{1},i_{2},\cdots,i_{K})\in\{1,2,\cdots,n\}^{K}:a_{i_{1}}^{1}\wedge a_{i_{2}}^{2}\wedge\cdots\wedge a_{i_{K}}^{K}>0\}.$
@@
$\mathcal{N}_{\phi(y_{k})}(\text{namely,}~P_{k}(I))=\{p_{k,l}:1\leq l\leq M_{k}\},$
@@
that is to say, $I_{k,l}=P_{k}^{-1}(p_{k,l})$. Then it is clear that $I=\bigcup_{l=1}^{M_{k}}I_{k,l}$ and $\{I_{k,l}:1\leq l\leq M_{k}\}$ is a pairwise disjoint family. Now, by (2) as has been proved, there exists a family $\{\varphi_{k,l}:1\leq l\leq M_{k}\}$ of proper labeling functions of $\mathscr{S}^{\prime}$ such that $\varphi_{k,l}(y_{k})=p_{k,l}$ for each $l\in\{1,2,\cdots M_{k}\}$, namely,
@@
For each $k\in\{1,2,\cdots,K\}$ and each $l\in\{1,2,\cdots,M_{k}\}$, we define a proper labeling function $\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}$ of $\mathscr{S}^{\prime}$ for each $(i_{1},i_{2},\cdots,i_{K})\in I_{k,l}$ by $\psi_{(i_{1},i_{2},\cdots,i_{K})}^{k}=\varphi_{k,l}$. Then one can easily see that
@@
Lemma 3.3, Lemma 3.4 and Theorem 3.5 have shown that $L^{0}$-simplexes, $L^{0}$-simplicial subdivisions and proper $L^{0}$-labeling functions are closely connected to their classical counterparts. These results motivate us to make full use of Lemma 2.8 to complete the proof of Theorem 2.9.
We are now in a position to prove Theorem 2.9.
@@
Further, let $S_{m}=Conv_{L^{0}}(\{y_{1}^{m},\cdots,y_{n}^{m}\})$ for each $m\in\{1,2,\cdots,M\}$. Then by part (1) of Lemma 3.3,
@@
is an (n1)(n-1)-dimensional $L^{0}$-simplex contained in $S$. Furthermore, by (3.1), we have
@@
4 Proof of Theorem 2.1: The Random Brouwer Fixed Point Theorem
@@
Definition 4.1 here is adopted from [20, 21] by following the traditional nomenclature of random metric spaces and random normed spaces (see [50, Chapters 9 and 15]).
@@
An ordered pair $(E,\|\cdot\|)$ is called a random normed module (briefly, an $RN$ module) over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$ if $E$ is an $L^{0}$-module and $\|\cdot\|$ is a mapping from $E$ to $L^{0}_{+}(\mathcal{F})$ such that the following conditions are satisfied:
@@
Clearly, when $(\Omega,\mathcal{F},P)$ is trivial, namely, $\mathcal{F}=\{\Omega,\emptyset\}$, then an $RN$ module with base $(\Omega,\mathcal{F},P)$ just reduces to an ordinary normed space.
Let $(E,\|\cdot\|)$ be an $RN$ module $(E,\|\cdot\|)$ over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$. For any real numbers $\varepsilon>0$ and $0<\lambda<1$, let
@@
Then $\mathcal{U}_{\theta}=\{U_{\theta}(\varepsilon,\lambda):\varepsilon>0,~0<\lambda<1\}$ forms a local base of some metrizable linear topology for $E$, called the $(\varepsilon,\lambda)$-topology induced by $\|\cdot\|$, denoted by $\mathcal{T}_{\varepsilon,\lambda}$. The idea of introducing the $(\varepsilon,\lambda)$-topology for $RN$ modules originates from Schweizer and Sklar’s work on the theory of probabilistic metric spaces [50]. It is clear that a sequence $\{x_{n},n\in\mathbb{N}\}$ in $E$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $x\in E$ iff $\{\|x_{n}-x\|,n\in\mathbb{N}\}$ converges in probability $P$ to $\theta$. On the other hand, for any $x\in E$ and any $r\in L_{++}^{0}(\mathcal{F})$, let
$B(x,r)=\{y\in E:\|x-y\|<r~\text{on}~\Omega\}.$
Then $\{B(x,r):x\in E,r\in L_{++}^{0}(\mathcal{F})\}$ forms a base for some Hausdorff topology on $E$, which amounts to the locally $L^{0}$-convex topology [16] induced by $\|\cdot\|$, denoted by $\mathcal{T}_{c}$. It is well known from [24] that $(E,\mathcal{T}_{\varepsilon,\lambda})$ is a metrizable topological module over the topological algebra $(L^{0}(\mathcal{F},\mathbb{K}),\mathcal{T}_{\varepsilon,\lambda})$, and it is also well known from [16] that $(E,\mathcal{T}_{c})$ is a topological module over the topological ring $(L^{0}(\mathcal{F},\mathbb{K}),\mathcal{T}_{c})$.
@@
Define the $L^{0}$-inner product $\langle\cdot,\cdot\rangle:L^{0}(\mathcal{F},\mathbb{K}^{d})\times L^{0}(\mathcal{F},\mathbb{K}^{d})\rightarrow L^{0}(\mathcal{F},\mathbb{K})$ by
@@
where $\bar{y}_{i}$ denotes the complex conjugate of $y_{i}$ for each $i\in\{1,2,\cdots,n\}$. Then $(L^{0}(\mathcal{F},\mathbb{K}^{d}),\langle\cdot,\cdot\rangle)$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete random inner product module (briefly, a $\mathcal{T}_{\varepsilon,\lambda}$-complete $RIP$ module, see [23] for details). let $\|x\|=\sqrt{\langle x,x\rangle}$ for any $x\in L^{0}(\mathcal{F},\mathbb{K}^{d})$. Then $(L^{0}(\mathcal{F},\mathbb{K}^{d}),\|\cdot\|)$ is a $\mathcal{T}_{\varepsilon,\lambda}$-complete $RN$ module.
@@
The randomized version of the Bolzano-Weierstrass theorem [32] states that for any a.s. bounded sequence $\{\xi^{0}_{n},n\in\mathbb{N}\}$ of random vectors from $(\Omega,\mathcal{F},P)$ to $\mathbb{R}^{d}$ (namely, $\sup\{|\xi^{0}_{n}(\omega)|:n\in\mathbb{N}\}<+\infty$ $a.s.$) there exists a sequence $\{n_{k}^{0},k\in\mathbb{N}\}$ of positive integer-valued random variables on $(\Omega,\mathcal{F},P)$ such that $\{n_{k}^{0}(\omega),k\in\mathbb{N}\}$ is strictly increasing for each $\omega\in\Omega$ and $\{\xi^{0}_{n_{k}^{0}},k\in\mathbb{N}\}$ converges a.s., where $\xi^{0}_{n_{k}^{0}}(\omega)=\xi^{0}_{n_{k}^{0}(\omega)}(\omega)=\sum^{\infty}_{l=1}I_{(n_{k}^{0}=l)}(\omega)\xi^{0}_{l}(\omega)$ for any $k\in\mathbb{N}$ and any $\omega\in\Omega$ (as usual, $(n_{k}^{0}=l)$ denotes the set $\{\omega\in\Omega:n_{k}^{0}(\omega)=l\}$ for any $k$ and $l$ in $\mathbb{N}$). Motivated by this basic result, Guo et al. [27] introduced the following notion of random sequential compactness in $RN$ modules.
@@
(1) In Definition 4.3, we require $n_{k}$ to be an element in $L^{0}(\mathcal{F},\mathbb{N})$, rather than a positive integer-valued random variable as in [27, Definition 2.10]. This choice is made for simplifying written expressions, and it is easy to check that the two formulations are essentially equivalent. (2) In the sense of Definition 4.3, the randomized Bolzano-Weierstrass theorem can be restated as: any a.s. bounded nonempty subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is random relatively sequentially compact. Moreover, as pointed out in [27], if $G$ is $\sigma$-stable, then $G$ is random relatively sequentially compact iff $\bar{G}_{\varepsilon,\lambda}$ (the closure of $G$ under $\mathcal{T}_{\varepsilon,\lambda}$) is random sequentially compact. Consequently, any a.s. bounded $\sigma$-stable nonempty $\mathcal{T}_{\varepsilon,\lambda}$-closed subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (for example, an $L^{0}$-simplex) must be random sequentially compact. (3) For any (n1)(n-1)-dimensional $L^{0}$-simplex $S$ of an $RN$ module, as pointed out by [27, Lemma 5.5], there exists an $L^{0}$-affine $\mathcal{T}_{\varepsilon,\lambda}$-isomorphism $T$ from $\Delta_{n-1}$ to $S$, where $\Delta_{n-1}:=Conv_{L^{0}}(\{e_{1},e_{2},\cdots,e_{n}\})$ is the standard (n1)(n-1)-dimensional $L^{0}$-simplex of $L^{0}(\mathcal{F},\mathbb{R}^{n})$. Since $\Delta_{n-1}$ is random sequentially compact as discussed above, and $T$ is $\sigma$-stable and $\mathcal{T}_{\varepsilon,\lambda}$-continuous, then $S$ is random sequentially compact.
@@
    $a.s.$ sequentially continuous at $x_{0}\in G_{1}$ if $\{\|T(x_{n})-T(x_{0})\|_{2},n\in\mathbb{N}\}$ converges $a.s.$ to 0 for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G_{1}$ such that $\{\|x_{n}-x_{0}\|,n\in\mathbb{N}\}$ converges $a.s.$ to 0. Further, $T$ is said to be $a.s.$ sequentially continuous if $T$ is $a.s.$ sequentially continuous at any point in $G_{1}$.
@@
It is clear that a.s. sequential continuity implies $\mathcal{T}_{\varepsilon,\lambda}$-continuity. Moreover, when $G_{1}$ is $\sigma$-stable, the $\mathcal{T}_{\varepsilon,\lambda}$-continuity of a mapping $T$ from $G_{1}$ to $G_{2}$ necessarily implies random sequential continuity of $T$. Furthermore, [27, Lemma 4.3] shows that a $\sigma$-stable mapping $T$ of $G_{1}$ to $G_{2}$ is random sequentially continuous iff it is $\mathcal{T}_{c}$-continuous.
For any (n1)(n-1)-dimensional $L^{0}$-simplex $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ in an $RN$ module $E$ and any $L^{0}$-simplicial subdivision $\mathscr{S}=\{S_{i}:i\in I\}$ of $S$, the random diameter of $\mathscr{S}$ is defined by
@@
Let $(E,\|\cdot\|)$ be an $RN$ module over $\mathbb{K}$ with base $(\Omega,\mathcal{F},P)$, $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ an (n1)(n-1)-dimensional $L^{0}$-simplex in $E$ and $\mathscr{S}$ an $L^{0}$-barycentric subdivision of $S$. Then
@@
First, we claim that $diam(G)=\bigvee_{k,l=1}^{m}\|y_{i}-y_{j}\|$ for any (n1)(n-1)-dimensional $L^{0}$-simplex $G=Conv_{L^{0}}(\{y_{1},y_{2},\cdots,y_{n}\})$ in $E$. In fact, it is clear that $\bigvee_{k,l=1}^{m}\|y_{k}-y_{l}\|\leq diam(G)$. On the other hand, for any $x=\sum_{i=1}^{m}\lambda_{i}y_{i}$ and $y=\sum_{i=1}^{m}\mu_{i}y_{i}$ in $G$, where $\lambda_{i},\mu_{i}\in L_{+}^{0}(\mathcal{F})$ for any $i\in\{1,2,\cdots,m\}$ and $\sum_{i=1}^{m}\lambda_{i}=\sum_{i=1}^{m}\mu_{i}=1$, we have
@@
Second, for any $\pi\in G(n)$ and any $k,l\in\{1,2,\cdots,n\}$, without loss of generality, we can assume that $k>l$. Then
@@
The key idea of the proof for the classical Brouwer fixed point theorem for a continuous self-mapping on a simplex $S^{\prime}$ lies in constructing a proper labeling function $\psi$ on a given simplicial subdivision $\mathscr{S}^{\prime}$ of $S^{\prime}$ that is compatible with the continuous self-mapping on $S^{\prime}$, see [1, Theorem 6.1] for details. Lemma 4.8 below provides a method for constructing such a proper $L^{0}$-labeling function in connection with an (n1)(n-1)-dimensional $L^{0}$-simplex $S$ in an $RN$ module that improves the formulation of [12, Lemma 2.2] in that the single proper $L^{0}$-labeling function provided here can be applied to the $L^{0}$-simplicial subdivision of every (n1)(n-1)-dimensional $L^{0}$-simplex contained in $S$.
@@
Let $(E,\|\cdot\|)$ be an $RN$ module with base $(\Omega,\mathcal{F},P)$, $S=Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ an (n1)(n-1)-dimensional $L^{0}$-simplex in $E$ and $f:S\rightarrow S$ be a $\sigma$-stable mapping. Then there exists a mapping $\phi:S\rightarrow L^{0}(\mathcal{F},\{1,2,\cdots,n\})$ such that the following conditions are satisfied: